OpenStudy (bookworm14):

Does anyone have the time to check about 5-7 problems for me? I already have my answers. This is new material (sin/cos/tan) and I would like to make sure I am doing it right. Thank you ! I will post the first 2 questions below!

OpenStudy (bookworm14):

I circled my answers

OpenStudy (danjs):

good job

OpenStudy (goformit100):

I hope you haven't solve these questions using calculator.

OpenStudy (danjs):

10*cos(36) 18*sin(35)

OpenStudy (bookworm14):

Did i get it wrong?! :(

OpenStudy (danjs):

look good to me, just practice setting up the cos(a) and sin(a) ratios

OpenStudy (goformit100):

You aren't wrong.

OpenStudy (danjs):

tangent uses the sides... inverse tangent of that ratio ..

OpenStudy (danjs):

and use cosine for the last

OpenStudy (bookworm14):

Dan, is that for the next set i put (26 & 27) or the first set (21-22)?

OpenStudy (danjs):

all look good... what are you unsure about?

OpenStudy (bookworm14):

Trigonometry is just new to me, so I am having you all check my answers to make sure my mind is on the right track how about this, I got B for #28 >

OpenStudy (danjs):

ok, for 28, what did you do

OpenStudy (danjs):

you can just type the thought process if you want

OpenStudy (bookworm14):

For 28, I am given the hypotenuse and opposite so I used sin. method 8/14 (opp/hyp) = 0.571 the I did sin^-1 (0.571)= 34.8 (and then I have to round up to 35)

OpenStudy (danjs):

right, after a few practice, using sin, cos, tan, just becomes like a pattern, look opposite side and hyp for sin, not really thinking opposite and hypotenuse...

OpenStudy (bookworm14):

29. B 30. B these are the last 2 I need check, you know, just for assuring purposes lol

OpenStudy (bookworm14):

Sorry I had so many questinos, I am just very prone to making simple mistakes and getting things wrong, even when I knw how to do it xD

OpenStudy (danjs):

inverse tangent, and inverse sin of those ratios

OpenStudy (danjs):

here, some examples

OpenStudy (danjs):

here is a reference sheet for things

OpenStudy (bookworm14):

Alrighty Thank you for checking them!! do you think you have time to walk me through one problem that I do not know how to work out that involves manipulating a formula? (it is the only one I cannot figure out on my whole assignment)

OpenStudy (danjs):

sure

OpenStudy (bookworm14):

alright, i am pretty good at formulas, but this one i have no idea

OpenStudy (danjs):

ok, that is just some algebra , moving things around

OpenStudy (bookworm14):

im gonna right down the steps as we go if thats all right (for future reference0

OpenStudy (danjs):

\[\large A = 2\pi*r^2 + 2\pi*r*h\]

OpenStudy (danjs):

ok,..notice that is the 2* area of a circle + circumference*height, or the surface area of a cylinder you want so solve for pi = something

OpenStudy (bookworm14):

I have to end up with an equation that solves for \[\pi = ?\] not numerical values, but variables

OpenStudy (danjs):

2 steps is easiest, factor the common pi from both of those terms \[\large A = 2\pi*r^2 + 2\pi*r*h = \pi*[2r^2 + 2rh]\]

OpenStudy (danjs):

then just divide by all that in the quantity

OpenStudy (danjs):

left with pi = A/[ ]

OpenStudy (bookworm14):

\[\pi = \frac{ a }{ 2r^2 + 2rh }\] ?

OpenStudy (danjs):

yep, factor the pi out, then divide by all the [ ] stuff, leaving pi by itself

OpenStudy (bookworm14):

ooh ok, see I was trying to factor 2pi r or something like that, idk what I was doing, but your way makes perfect sense!

OpenStudy (bookworm14):

Thank you so much for taking the time to help me! I really feel like I learned! Thank you!

OpenStudy (danjs):

yes, the 2 and the r, are common factors there, but in the end you want pi by itself, just more work your way no prob, anytime